Suppose a pie factory’s production function is Q=2KLQ=2KL, where Q is the quantity of pies produced...

  1. Suppose a pie factory’s production function is Q=2KLQ=2KL, where Q is the quantity of pies produced each hour, K is the number of ovens used, and L is the number of workers per hour.
    1. If KK is fixed at 5 units, what is the firm’s total product of labor equation (i.e., its short-run production function)?
    2. Assume 3 units of labor are employed with the 5 units of capital. If production achieves technological efficiency, how many pies are produced per hour?
    3. If 4 workers are employed, and 35 pies are produced, what can be said about efficiency?
Answer #1

I=2KL Put K = 5 @= 2(5) L - 0=102 (Shand run production and O = 2KL put 1-5 L = 3 = 2(5) (3) 10 = 30 with 3 units of labor ,